# CubistMerge: Spatial-Preserving Token Merging For Diverse ViT Backbones

### v1

@@ Abstract
Many modern ViT backbones adopt spatial architectural designs, such as window attention, decomposed relative positional embeddings in SAM, and RoPE in DINOv3. Such architectures impose new challenges on token reduction, as the vast majority of existing methods fail to preserve the spatial structure these architectures depend on.
In this paper, we introduce a simple yet effective token merging method that maintains spatial integrity, enabling seamless compatibility with spatial architectures. We reconcile two seemingly conflicting requirements: (i) exploiting the uneven information distribution across the spatial layout while (ii) preserving the spatial structure post-merging. Our approach employs (i) a 2D reduction strategy to enforce structured token layouts, (ii) a spatial-aware merging algorithm that maintains relative token positions, and (iii) a novel max-magnitude-per-dimension token representation that preserves salient features.
Our method demonstrates strong performance both off-the-shelf and with fine-tuning, achieving state-of-the-art results on spatial and non-spatial architectures across various vision tasks. Specifically, we achieve 1.25× speedup on SAM-H with only 0.7% mIOU drop evaluated on COCO off-the-shelf, and 1.15× speedup on DeiT-B with no top-1 accuracy drop on ImageNet within just one epoch of fine-tuning.
@@ 1 Introduction
Token reduction methods offer an attractive solution by leveraging the input-agnostic nature of transformers to dynamically reduce the number of tokens during processing. However, the vast majority of existing token reduction methods face fundamental incompatibilities with spatial architectures, such as 2D positional embeddings (Li et al., 2022b; Heo et al., 2024) at every attention layer, and window attention (Liu et al., 2021; Li et al., 2022a). Most techniques (Rao et al., 2021; Liang et al., 2022b; Kong et al., 2022; Bolya et al., 2023; Lee et al., 2024a; Tran et al., 2024; Long et al., 2023; Kim et al., 2024; Norouzi et al., 2024; Chen et al., 2023) produce unstructured token layouts that break spatial coherence (see Figure 1). The resulting unstructured token layouts break both window attention, which requires consistent token counts across all windows, as well as 2D positional embeddings, which depend on structured arrangements to compute spatial relationships correctly.
@@ 2 Related Work
Retraining-Based Token Reduction. Some token reduction approaches require extensive retraining (Long et al., 2023; Xu et al., 2024; Lu et al., 2023; Liu et al., 2024b; Lee and Hong, 2024; Liu et al., 2024a), which presents challenges for modern large-scale models due to computational costs of training and limited data availability: foundation models such as DINOv3 (Siméoni et al., 2025) rely on massive datasets and scale architectures up to 7B parameters. While effective, these retraining-based approaches are prohibitively expensive, creating a need for training-free solutions.
Graph-based Token Merging. Token Merging (ToMe) (Bolya et al., 2023) addresses both limitations above: it merges tokens rather than discarding them, enabling recovery for dense outputs, and can operate off-the-shelf without retraining. ToMe employs a graph-based approach with bipartite matching to selectively combine similar tokens through weighted averaging. This approach demonstrates success across several models and tasks, becoming the foundation for subsequent works with incremental improvements such as adaptive merging rates (Norouzi et al., 2024; Chen et al., 2023), importance-based token selection (Lee et al., 2024a; Tran et al., 2024; Long et al., 2023), and hybrid pruning-merging approaches (Kim et al., 2024). However, ToMe and these subsequent works all fail to maintain spatial structure after merging, which is critical for architectures with spatial components. Despite some works adopting spatial-aware strategies (Norouzi et al., 2024), they only focus on merging spatially near tokens but do not maintain structured spatial layouts in resulting tokens.
@@ 3 Methods
Both phases use our spatial-aware token matching algorithm described in Section 3.2.
Figure 2: 2D token reduction with spatial-aware merging: (1) original 14×14 tokens, (2) select horizontal tokens to merge, (3) merge horizontally to 14×12 tokens, (4) select vertical tokens to merge, (5) merge vertically to 12×12 tokens.
@@ 3.2 Spatial-Aware Token Matching
Graph Construction. Since our 2D reduction operates on each row and column independently, tokens within each subset naturally form a linear arrangement based on their spatial positions. Motivated by this, we use a path graph to further preserve spatial coherence, where each token only connects to its adjacent neighbors within the same row or column. This design ensures that merged tokens maintains the original relative spatial positions of their constituent tokens, which is critical for 2D positional embeddings (Li et al., 2022b; Heo et al., 2024).
@@
To enable better parallelization, we adopt ToMe’s (Bolya et al., 2023) node bipartition approach, which alternates token role assignments so that adjacent tokens have complementary roles (source and destination). Each source token then nominates its most similar adjacent neighbor as its merge destination, and we select the top-k edges from these nominations. This guarantees that no more than three tokens are ever merged, so merging never takes more than two steps. However this does not strictly guarantee the selection of the most similar edges.
@@ 4.1 Spatial Architectures
4.1.1 Model Sweep
MViTv2. We evaluate MViTv2 (Li et al., 2022b), which uses decomposed relative positional embeddings, on image classification. Table 4.1 presents layer sensitivity results and includes fine-tuning results for MViTv2-B, with fine-tuning limited to 3 epochs. For meaningful comparisons, we use $l=10$ for MViTv2-B and $l=20$ for MViTv2-L, where Expedite and ToMe achieve more reasonable performance, and vary $r_{h}$, $r_{w}$ to produce the results shown in Figure 3.
DINOv3. We evaluate image classification and object detection using ViT7B backbone (Siméoni et al., 2025), which incorporates RoPE for 2D positional embeddings. Classification results are shown in Table 4.1 and Figure 3, while object detection results are shown in Table 4 and Figure 4. We experimented at $l=10$ and $l=20$, and selected $l=20$ where Expedite and ToMe exhibits more reasonable results, for further experiments varying $r_{h}$ and $r_{w}$. Despite using the same pretrained backbone, ToMe and Expedite exhibit much worse layer sensitivity for object detection at $l=10$, while CubistMerge maintains consistent performance across both tasks.
@@ 4.2 Non-Spatial Architectures
DeiT We evaluate on DeiT-B compared against ToMe (Bolya et al., 2023), PiToMe (Tran et al., 2024), ToFu (Kim et al., 2024), MCTF (Lee et al., 2024a) and DynamicViT (Rao et al., 2021). We evaluate both off-the-shelf and with fine-tuning limited to 5 epochs. Figure 8 shows results using existing methods’ recommended merging schedules and CuMe applied at $l=1$. Fine-tuned results are shown in Table 8. CuMe achieves state-of-the-art results with no accuracy loss at 1.15× speedup within just one epoch of fine-tuning, while maintaining competitive performance across higher speedup ratios.

### v2

# arXiv:2509.21764v2[cs.CV] 02 Mar 2026

@@ Abstract
We introduce a simple yet effective token merging method for ViTs that is compatible with modern spatial ViT architectures like SAM or DINOv3, by maintaining spatial integrity of merged tokens. Our proposal reconciles two seemingly conflicting requirements: (i) exploiting the uneven information distribution across the spatial layout while (ii) preserving the spatial structure post-merging. Our approach employs (i) a 2D reduction strategy to enforce structured token layouts, (ii) a spatial-aware merging algorithm that maintains relative token positions, and (iii) a novel max-magnitude-per-dimension token representation that preserves salient features. Our method demonstrates strong performance both off-the-shelf and with fine-tuning, achieving state-of-the-art results on spatial and non-spatial architectures across various vision tasks. Specifically, we achieve 1.25× speedup on SAM-H with only 0.7% mIOU drop evaluated on COCO off-the-shelf, and 1.15× speedup on DeiT-B with no top-1 accuracy drop on ImageNet within just one epoch of fine-tuning.
@@ 1 Introduction
Token reduction methods offer an attractive solution by leveraging the input-agnostic nature of transformers to dynamically reduce the number of tokens during processing. However, the vast majority of existing token reduction methods face fundamental incompatibilities with spatial architectures, such as 2D positional embeddings (Li et al., 2022b; Heo et al., 2024) at every attention layer, and window attention (Liu et al., 2021; Li et al., 2022a). Most techniques (Rao et al., 2021; Liang et al., 2022b; Kong et al., 2022; Bolya et al., 2023; Lee et al., 2024a; Tran et al., 2024; Long et al., 2023; Kim et al., 2024; Norouzi et al., 2024; Chen et al., 2023) produce unstructured token layouts that break spatial coherence (see Figure 1). The resulting unstructured token layouts break both window attention, which requires consistent token counts across all windows, as well as 2D positional embeddings, which depend on structured arrangements to compute spatial relationships correctly. The impact of breaking spatial coherence is shown in Figure 2(b): non-spatial-preserving methods like ToMe severely distort attention patterns of models with relative positional bias. In contrast, our spatial-preserving approach maintain attention patterns that closely resemble the baseline model, as shown in Figure 2(a). [^1] See Appendix A.1 for a more in-depth background on why spatial coherence is critical for modern architectures.
@@
(a) Original
(b) ToMe
(c) Ours
Figure 2: Attention patterns with relative positional bias towards 5 different token positions (indicated by red star) on SAM-B. (a) shows attention map of baseline model. (b) shows effective attention pattern with ToMe applied (c) shows effective attention pattern with our method applied. Our method preserves attention patterns better than non-spatial-preserving method like ToMe.
@@ 2 Related Work
Retraining-Based Token Reduction. Some token reduction approaches require extensive retraining (Long et al., 2023; Lu et al., 2023; Liu et al., 2024b; Lee and Hong, 2024; Liu et al., 2024a), which presents challenges for modern large-scale models due to computational costs of training and limited data availability: foundation models such as DINOv3 (Siméoni et al., 2025) rely on massive datasets and scale architectures up to 7B parameters. While effective, these retraining-based approaches are prohibitively expensive, creating a need for training-free solutions.
Graph-based Token Merging. Token Merging (ToMe) (Bolya et al., 2023) addresses both limitations above: it merges tokens rather than discarding them, enabling recovery for dense outputs, and can operate off-the-shelf without retraining. ToMe employs a graph-based approach with bipartite matching to selectively combine similar tokens through weighted averaging. This approach demonstrates success across several models and tasks, becoming the foundation for subsequent works with incremental improvements such as adaptive merging rates (Norouzi et al., 2024; Chen et al., 2023), importance-based token selection (Lee et al., 2024a; Tran et al., 2024; Long et al., 2023), and hybrid pruning-merging approaches (Kim et al., 2024). However, ToMe and these subsequent works all fail to maintain spatial structure after merging, which is critical for architectures with spatial components. Despite some works adopting spatial-aware strategies (Norouzi et al., 2024; Xu et al., 2024), they only focus on merging spatially near tokens but do not maintain structured spatial layouts in resulting tokens.
@@ 3 Methods
Both phases use our spatial-aware token matching algorithm described in Section 3.2. Algorithm 1 gives the full tensor-level implementation. [^2] For dense prediction tasks requiring full spatial resolution, the 2D reduction is invertible. Appendix A.3 details the recovery procedure.
Algorithm 1: 2D Token Reduction Input: $X\in\mathbb{R}^{N\times H\times W\times D}$ Output: $X^{\prime\prime}\in\mathbb{R}^{N\times(H-r_{h})\times(W-r_{w})\times D}$ // Horizontal Reduction $X=X.\text{transform}([N,H,W,D]\rightarrow[N\times H,W,D])$ $X^{\prime}=\textsc{BatchTokenMerge}(X,r_{w})$ // Vertical Reduction $X^{\prime}=X^{\prime}.\text{transform}([N\times H,(W-r_{w}),D]\rightarrow[N\times(W-r_{w}),H,D])$ $X^{\prime\prime}=\textsc{BatchTokenMerge}(X^{\prime},r_{h})$ return $X^{\prime\prime}.\text{transform}([N\times(W-r_{w}),(H-r_{h}),D]\rightarrow[N,(H-r_{h}),(W-r_{w}),D])$
Figure 3: 2D token reduction with spatial-aware merging: (1) original 14×14 tokens, (2) select horizontal tokens to merge, (3) merge horizontally to 14×12 tokens, (4) select vertical tokens to merge, (5) merge vertically to 12×12 tokens.44footnotemark: 4
@@ 3.2 Spatial-Aware Token Matching
Graph Construction. Since our 2D reduction operates on each row and column independently, tokens within each subset naturally form a linear arrangement based on their spatial positions. Motivated by this, we use a path graph to further preserve spatial coherence, where each token only connects to its adjacent neighbors within the same row or column. This design ensures that merged tokens maintain the original relative spatial positions of their constituent tokens, which is critical for 2D positional embeddings (Li et al., 2022b; Heo et al., 2024). Additionally, path graphs minimize the number of edges by connecting only adjacent tokens, reducing the complexity of computing similarity scores from $O(N^{2})$ to $O(N)$ compared to global token merging algorithms such as ToMe (Bolya et al., 2023) that computes all pairwise token similarities.
Our path graph construction enforces a strict adjacency constraint: tokens can only merge with their immediate spatial neighbors. This guarantees that the merged token’s position maintains the original spatial ordering of its constituent tokens. In contrast, global matching approaches like ToMe (Bolya et al., 2023) merging across the entire token sequence: a token at position 0 may merge with one at position W1W-1, with the resulting merged token placed at $(W-r_{w})-1$. Such long-range merging fundamentally disrupts the spatial correspondence between token positions and their original locations. While some spatially-aware methods like GTP-ViT (Xu et al., 2024) incorporate spatial proximity as a factor of consideration, they do not enforce adjacency as a hard constraint. Consequently, these methods still permit merging between spatially distant tokens when similarity scores favor such pairings, failing to guarantee preservation of relative spatial relationship in the output token arrangement.
@@
To enable better parallelization, we adopt ToMe’s (Bolya et al., 2023) node bipartition approach, which alternates token role assignments so that adjacent tokens have complementary roles (source and destination). Each source token then nominates its most similar adjacent neighbor as its merge destination, and we select the top-k edges from these nominations (see Figure 4(c)). This guarantees that no more than three tokens are ever merged, so merging never takes more than two steps. However this does not strictly guarantee the selection of the most similar edges. [^5] See Appendix A.2 for a thorough analysis of design choices, including naive selection, reduction-tree optimization, and bipartite matching.
(a) naive,local
(b) naive,local ($\log N$)
(c) bipartite,local (ours)
Figure 4: Illustration of edge selection algorithms. Arrows on selected edges (in orange) indicate the direction of token merging, pointing from the source token to the destination token. The numbers on edges represent execution order of merging required by dependencies. (a) Path graph with naive edge selection, requiring sequential execution. (b) Path graph with naive edge selection, optimized with reduction tree to $O(\log N)$ complexity. (c) Path graph with bipartite edge selection to eliminate dependencies by ensuring each source token (in red) can only merge to one destination (in blue).
@@ 4.1 Spatial Architectures
4.1.1 Model Sweep
MViTv2. We evaluate MViTv2 (Li et al., 2022b), which uses decomposed relative positional embeddings, on image classification. Table 4.1.1 presents layer sensitivity results and includes fine-tuning results for MViTv2-B, with fine-tuning limited to 3 epochs. For meaningful comparisons, we use $l=10$ for MViTv2-B and $l=20$ for MViTv2-L, where Expedite and ToMe achieve more reasonable performance, and vary $r_{h}$, $r_{w}$ to produce the results shown in Figure 5.
DINOv3. We evaluate image classification and object detection using ViT7B backbone (Siméoni et al., 2025), which incorporates RoPE for 2D positional embeddings. Classification results are shown in Table 4.1.1 and Figure 5, while object detection results are shown in Table 6 and Figure 6. We experimented at $l=10$ and $l=20$, and selected $l=20$ where Expedite and ToMe exhibits more reasonable results, for further experiments varying $r_{h}$ and $r_{w}$. Despite using the same pretrained backbone, ToMe and Expedite exhibit much worse layer sensitivity for object detection at $l=10$, while CubistMerge maintains consistent performance across both tasks.
@@ 4.2 Non-Spatial Architectures
DeiT We evaluate on DeiT-B compared against ToMe (Bolya et al., 2023), PiToMe (Tran et al., 2024), ToFu (Kim et al., 2024), MCTF (Lee et al., 2024a), GTP-ViT (Xu et al., 2024) and DynamicViT (Rao et al., 2021). We evaluate both off-the-shelf and with fine-tuning limited to 5 epochs. Figure 10 shows results using existing methods’ recommended merging schedules and CuMe applied at $l=1$. Fine-tuned results are shown in Table 10. CuMe achieves state-of-the-art results with no accuracy loss at 1.15× speedup within just one epoch of fine-tuning, while maintaining competitive performance across higher speedup ratios.
@@
4.3 Discussion
Our experiments demonstrate the critical importance of preserving spatial structure in spatial architectures. ToMe’s failure to maintain spatial layouts causes severe performance degradation across experiments on spatial architectures in Section 4.1: on DINOv3-ViT7B at $l=10$, ToMe achieves only 84.3% top-1 versus CuMe’s 87.7% (Table 4.1.1) for classification, and catastrophic 31.0 AP versus CuMe’s 55.5 on detection tasks (Table 6). As visualized in Figure 2, methods like ToMe distort attention patterns by disrupting spatial coherence in merged tokens.
While Expedite preserves spatial structure, its uniform pooling-based approach fails to exploit uneven information distribution. This manifests in poor layer sensitivity: Expedite suffers severe metric drops at early layers on SAM, ViTDet, and Mask2Former, requiring careful layer selection to achieve reasonable performance. In contrast, CuMe’s selective merging maintains consistent performance across layers by preserving distinct tokens while merging only redundant ones. CuMe outperforms Expedite across every spatial architecture experiment in Section 4.1.
@@ 5 Conclusion
Appendix A Appendix
A.1 Challenges of Spatial Architectures
A.1.1 Window Attention
Background. The Swin Transformer (Liu et al., 2021) introduced sliding window attention to address the quadratic complexity of global self-attention in vanilla ViT (Dosovitskiy et al., 2021). By restricting self-attention to non-overlapping local windows, Swin achieves linear complexity while maintaining modeling capacity through shifted windowing for cross-window connections (see Figure 11). This mechanism has since become foundational in subsequent architectures. ViTDet (Li et al., 2022a) validated the effectiveness of window attention for dense prediction tasks, and showed that simpler window attention without shifting is sufficient when aided by a few cross-window propagation blocks. This non-shifting variant was then adopted by state-of-the-art models (Kirillov et al., 2023; Ravi et al., 2025; Ryali et al., 2023).
Why Preserving Spatial Layout is Critical. Window attention operates on local spatial regions, leveraging the high correlation of nearby visual features (Liu et al., 2021). This requires tokens to maintain a coherent 2D spatial arrangement to be partitioned into windows, otherwise tokens may lose opportunities to attend to tokens in local regions, as shown in Figure 11’s demo of ToMe. The shifted windowing scheme further depends on this structured layout to enable cross-window connections.
Token reduction methods that fail to preserve spatial structure break this assumption: unstructured methods like ToMe (Bolya et al., 2023) produce irregular token layouts where different windows reduce varying numbers of tokens after merging. This leads to two unpalatable options: (1) maintaining the original window grouping — so that each window contains a different number of tokens — is at odds with the regular SIMD dataflow properties that accelerators like GPUs rely on for performance; while (2) padding all windows to the same length would offset the computation reduction benefit of token merging.
Another naive solution might be to reduce $H\times W$ tokens to $H^{\prime}\times W^{\prime}$ tokens and treat the reduced set as a new 2D token layout that can be partitioned into windows. This naive approach has also been used in LTM (Wang and Yang, 2025). However, as shown in Figure 11, this approach destroys the spatial correspondence between the original and reduced layouts, as windows may now group spatially distant tokens together while placing spatially local tokens in different windows, defeating the purpose of local window attention.
Figure 11: Window attention with shifted window partitioning. Cyan lines indicate window boundaries. Top row shows the original token layout with regular and shifted window partitioning. Middle and bottom rows demonstrate how token merging methods interact with window attention: ToMe (middle) destroys spatial structure, causing misalignment with window boundaries after merging, while CuMe (bottom) preserves the structured layout, maintaining compatibility with both regular and shifted window partitions.
A.1.2 2D Positional Embedding
Background. Decomposed relative positional embeddings were introduced in MViTv2 (Li et al., 2022b), encoding spatial relationships based on relative spatial distances between tokens. Unlike absolute positional embeddings in vanilla ViT  (Dosovitskiy et al., 2021), these learned parameters are injected into each attention layer and computed separately along height and width dimensions. SAM adopts this strategy, combining it with window attention for strong zero-shot capabilities (Kirillov et al., 2023). RoPE (Heo et al., 2024), on the other hand, encodes 2D spatial relationships through axial frequency operations applied, again, separately for x and y dimensions. DINOv3 (Siméoni et al., 2025) demonstrates RoPE’s effectiveness for self-supervised visual representation learning.
Figure 12: Visualization of relative positional embedding with attention heat map towards 5 different token positions indicated by red stars. Baseline (top): (a) positional bias component, (b) attention score without positional bias, and (c) attention score with positional bias (actual attention score). ToMe (middle): (d) attention score and (e) effective attention score on original spatial layout, recovered from attention score on merged layout. CuMe (bottom): (f) and (g) similarly for CuMe. Collected from block 2 of SAM-B.
Why Preserving Spatial Layout is Critical. 2D positional embeddings require tokens to maintain their relative spatial positions to correctly compute spatial relationships. When token reduction methods destroy the spatial structure, the positional embeddings can no longer accurately represent the spatial relationships between tokens, leading to significant performance degradation.
We demonstrate this with decomposed relative positional embeddings (Li et al., 2022b). As illustrated in Figure 12, positional bias (a) enhances attention from spatially near regions, as seen in the comparison between (b) and (c).
To understand the impact of token merging on spatial architectures that use positional embeddings, we visually compare attention patterns that appear in ToMe (Bolya et al., 2023) (non-spatial) and CuMe (spatial). For ToMe (d,e), we observe significant distortion from the baseline attention patterns, as the irregular token arrangement after merging disrupts the relative positional relationships that the embeddings depend on. In contrast, CuMe (f,g) maintains structured token layouts that preserve relative positions, and yields attention patterns that closely align with the baseline. This demonstrates why preserving spatial structure is critical for architectures relying on 2D positional embeddings.
A.2 Details on Graph Construction and Edge Selection
Figure 4 illustrates the graph construction and edge selection approaches discussed in Section 3.2.
Naive Edge Selection. Dependency chains arise when three or more adjacent tokens must be merged sequentially. For example, if edges b→c, c→d, and d→e are all selected, token c must first receive b before it can merge into d, forcing sequential execution with $O(N)$ time complexity in the worst case (Figure 4(a)).
Reduction-Tree-Optimized Naive Edge Selection. Reduction tree approaches can theoretically improve this to $O(\log N)$ by reorganizing the merge order (Figure 4(b)): tokens b→c and d→e merge simultaneously in the first step, then c→d is redirected to merge c→e in the second step. However, such tree-structured computations are not a good match for SIMD execution on GPUs and are less well supported by frameworks like PyTorch. Moreover, even with logarithmic complexity, the sequential dependencies still limit parallelization.
Bipartite Solution. The bipartite approach (Figure 4(c)) eliminates all inter-step dependencies, enabling fully parallel execution via scatter-reduce operations—avoiding the $O(N)$ and $O(\log N)$ sequential bottlenecks of the naive and reduction-tree alternatives respectively.
A.3 Token Recovery for Dense Prediction Tasks
A.3.1 2D Token Recovery
Token recovery is performed in reverse order of the reduction process: we first recover the vertical dimension, then the horizontal dimension, as illustrated in Figure 13.
A.3.2 Token Recovery within Row/Column
We employ the simple token recovery that has been implicitly used in ToMe (Bolya et al., 2023). Given multiple source tokens $\{x_{\text{src}_{1}},x_{\text{src}_{2}},\ldots,x_{\text{src}_{k}}\}\in\mathbb{R}^{c}$ that merge into a single destination token, we define:
$x_{\text{merged}}=\textsc{MergeTokens}([x_{\text{src}_{1}},x_{\text{src}_{2}},\ldots,x_{\text{src}_{k}}])$ (1)
For recovery, each source token is recovered by duplicating the merged token:
$x^{\prime}_{\text{src}_{i}}=x_{\text{merged}}\quad\forall i\in\{1,2,\ldots,k\}$ (2)
Note that our MergeTokens operation uses max-magnitude-per-dimension representation (see Section 3.3), which differs from the weighted averaging used in ToMe (Bolya et al., 2023).
(a) Reduced
12×12 tokens
(b) Vertical Recovery
14×12 tokens
(c) Horizontal Recovery
14×14 tokens
Figure 13: Token recovery process for CubistMerge. Starting from (a) the reduced 12×12 layout, we (b) first recover the vertical dimension to 14×12, then (c) recover the horizontal dimension to the original 14×14 layout. Grid lines show token boundaries at each stage.